# Correlated decoding of logical algorithms with transversal gates

### v1

@@ Abstract
Quantum error correction is believed to be essential for scalable quantum computation, but its implementation is challenging due to its considerable space-time overhead. Motivated by recent experiments demonstrating efficient manipulation of logical qubits using transversal gates (Bluvstein et al., Nature 626, 58-65 (2024)), we show that the performance of logical algorithms can be substantially improved by decoding the qubits jointly to account for physical error propagation during transversal entangling gates. We find that such correlated decoding improves the performance of both Clifford and non-Clifford transversal entangling gates, and explore two decoders offering different computational runtimes and accuracies. By considering deep logical Clifford circuits, we find that correlated decoding can significantly improve the space-time cost by reducing the number of rounds of noisy syndrome extraction per gate. These results demonstrate that correlated decoding provides a major advantage in early fault-tolerant computation, and indicate it has considerable potential to reduce the space-time cost in large-scale logical algorithms.
Quantum error correction (QEC) is believed to be essential for large-scale quantum computation Shor (1995); Preskill (1998); Dennis et al. (2002); Nielsen and Chuang (2010); Kitaev (2003). It is a rapidly developing frontier, with recent milestone demonstrations of the preservation of a logical qubit and universal logic gates Ofek et al. (2016); Reinhold et al. (2020); Ryan-Anderson et al. (2021); Postler et al. (2022); Ryan-Anderson et al. (2022); AI and Collaborators (2023a, b); Menendez et al. (2023); Sivak et al. (2023); Wang et al. (2023a); Self et al. (2024). However, implementing large-scale algorithms with protected, logical qubits is challenging due to its considerable space-time overhead. Most recently, reconfigurable neutral atom arrays have realized quantum algorithms with tens of logical qubits and hundreds of logical entangling gates Bluvstein et al. (2024), opening the door to early fault-tolerant computation. Key to these results was the ability to efficiently perform fault-tolerant transversal entangling gates by dynamically reconfiguring qubits during the computation Bluvstein et al. (2022), similar to prior work with trapped ions Postler et al. (2022); Ryan-Anderson et al. (2022); Menendez et al. (2023); Wang et al. (2023a).
@@
Correlated decoding algorithms.— To formalize this approach, we define the decoding hypergraph of the logical algorithm Gidney (2021); Gottesman (2022); Higgott et al. (2023); McEwen et al. (2023); Delfosse and Paetznick (2023), which relates each possible physical error mechanism to the stabilizer measurement(s) it flips. The hypergraph vertices correspond to $N$ measured checks of the logical qubits $\{C_{i}\}_{i=1\dots N}$, which compare consecutive stabilizer measurements in time. Each check is defined as the product of a stabilizer measurement and its previous measurement (if one exists), obtained from back-propagating the stabilizer operator through the circuit to the point it was previously measured. If one stabilizer measurement in the check detects an error, the check flips from $+1$ to $-1$. Each error mechanism is associated with a hyperedge connecting the check(s) it flips. Errors copied during transversal entangling gates therefore correspond to hyperedges connecting the checks of multiple logical qubits (Fig. 1(b)). We use Stim Gidney (2021), a Clifford circuit simulator, to identify the $M$ hyperedges $\{E_{j}\}_{j=1\dots M}$ and their probabilities $\{p_{j}\}_{j=1\dots M}$ under a chosen circuit error model, and which error source(s) $I(C_{i})$ flip each check $C_{i}$ (Fig. 1(c)).
@@
Because finding the MLE exactly is NP-hard, we expect the runtime of the above algorithm to grow exponentially in $M$ in the worst-case Berlekamp et al. (1978); Vardy (1997). This motivates our second algorithm, hypergraph union-find with belief propagation (belief-HUF) Delfosse and Nickerson (2021); Delfosse et al. (2022), which instead approximates the MLE and has an efficient time complexity of $O(M^{3}\log(M))$ sup . Belief-HUF first uses belief propagation to update the error probabilities $\{p_{j}\}_{j=1\dots M}$ to approximate the probability of each error given the measured syndrome Higgott et al. (2023). Leveraging the insight that errors are typically localized, belief-HUF then expands clusters of error mechanisms and checks on the decoding hypergraph at a rate dependent on the error probabilities, merging clusters that meet. A cluster stops expanding when it contains an error mechanism consistent with its measured checks, and the algorithm terminates when all clusters stop expanding.
Decoding transversal Clifford gates.— As a simple illustrative example, we apply these techniques to decode a transversal CNOT, which is available for all Calderbank-Shor-Steane (CSS) codes Shor (1997a). A key property of error-correcting circuits is that their performance can be improved for physical error rates below some threshold $p_{\text{th}}$ by increasing the code distance (the minimum weight of a logical Pauli operator), which is related to how many errors can be corrected Dennis et al. (2002); Fowler et al. (2012). The threshold fundamentally depends on the decoder, as more accurate decoders can tolerate higher physical error rates. To explore the optimal advantage of correlated decoding, we will compare decoding two surface codes jointly with MLE, and independently with MWPM Higgott (2021); Higgott and Gidney (2023).
@@
Decoding deep logical Clifford circuits.— In realistic quantum algorithms, stabilizer measurements are used to remove entropy from the logical qubit by tracking physical errors over time. A challenge, however, is that the stabilizer measurements are also noisy. As a result, entangling gates based on lattice surgery require $d$ rounds of syndrome extraction to ensure fault tolerance to these errors, creating considerable space-time overhead Horsman et al. (2012). We now explore if we can perform fewer than $d$ rounds between transversal CNOTs, by leveraging the deterministic propagation of stabilizer measurement errors through CNOTs to verify stabilizer measurements using surrounding rounds of syndrome extraction.
@@
In contrast, because the correlated decoders can be applied to general decoding hypergraphs, they can potentially decode circuits with fewer than $d$ syndrome extraction rounds between CNOTs. To study this, we numerically simulate deep logical Clifford circuits between four surface codes. The codes are initialized noiselessly in |$\ket{+_{L}}^{\otimes 4}$, then entangled with 32 layers of transversal gates, with $n_{r}$ rounds of syndrome extraction between layers (Fig. 3(b)). Each layer is comprised of a random transversal Hadamard $H_{L}$ or Pauli gate $\{X_{L},Y_{L},Z_{L}\}$, followed by a random pairing of CNOTs with randomized designation of control and target. We add depolarizing noise to each circuit-level physical operation with probability $p$, as in Refs. Higgott et al. (2023); sup . Finally, we noiselessly measure the $X$ and $Z$ stabilizers and the logical stabilizers of the resulting state, then decode to find the logical error rate $P_{L}$ Gottesman (1998). In the Supplementary Materials sup , we confirm that our results are insensitive to boundary effects from the final noiseless measurements, and representative of different randomly sampled circuits.
@@
We first study the case of one round of syndrome extraction per CNOT ($n_{r}=1$). Figure 3(c) shows the logical error rate per round, $P_{L,\text{max}}\big{[}1-(1-P_{L}/P_{L,\text{max}})^{1/(32n_{r})}\big{]}$, where $P_{L,\text{max}}=1-\frac{1}{2^{4}}$ is the error rate of the maximally mixed logical state, as a function of $p$ sup . The numerical results are consistent with the existence of a threshold for both belief-HUF and MLE. Though we cannot reliably fit the thresholds because $P_{L}/P_{L,\text{max}}\to 1$ near the threshold, we estimate that $p_{\text{th}}\geq 0.56\%$ for belief-HUF and $p_{\text{th}}\geq 0.87\%$ for MLE from the largest physical error rate at which $P_{L}$ decreases with $d$ sup . Furthermore, in the Supplementary Materials sup we observe that the MLE scaling is consistent with achieving the full code distance.
@@
Decoding transversal non-Clifford gates.— Correlated decoding can also be applied to non-Clifford transversal gates. As a proof-of-concept, we study a transversal CCZ between three fifteen-qubit quantum Hamming codes, with perfect syndrome extraction Paetznick and Reichardt (2013); sup . We perfectly initialize each code in an $X$ or $Z$-basis product state, then perform a CCZ with single-physical-qubit depolarizing noise beforehand. Because the CCZ is non-Clifford, these errors can propagate to a superposition of Pauli errors on multiple logical qubits, each generating a hyperedge in the decoding hypergraph (Fig. 5(a-b)) sup ; Scruby et al. (2022). We analyze initial states for which the circuit outputs a logical stabilizer state, which can be efficiently simulated sup . We then noiselessly measure the $X$ and $Z$ stabilizers and logical stabilizers, and decode to compute the average logical error rate over initial states Gottesman (1998). Figure 5(c) shows that correlated decoding with MLE outperforms decoding the qubits independently by a factor of $\simeq 1.5$. Note that the gain here is smaller than with Clifford gates, as the error propagation is no longer fully deterministic. This motivates additional analysis of correlated decoding with non-Clifford gates as an avenue for future research.
@@
Outlook.— Our results show that correlated decoding constitutes a powerful tool for improving the performance of logical algorithms. In particular, because stabilizer measurement errors can be tracked through the circuit, $d$ rounds of syndrome extraction are not necessarily required between transversal CNOTs. These insights were used in the experiments of Ref. Bluvstein et al. (2024) to improve the entanglement fidelity of a logical Bell pair between surface codes, and to fault-tolerantly prepare logical GHZ states. Looking forward, these theoretical results indicate that the number of syndrome extraction rounds can be optimized for each particular circuit. In circuits where $\simeq 1$ round per CNOT is optimal, the space-time cost can be reduced by a factor of $\simeq d$.
While this technique already offers key performance advantages in the era of early fault-tolerant quantum computation, in the long term it has the potential to substantially reduce the space-time cost of large-scale algorithms Shor (1997b); Lloyd (1996). In particular, correlated decoding can be potentially applied to core algorithmic subroutines, such as magic-state distillation Bravyi and Kitaev (2005); Fowler et al. (2012) and quantum arithmetic Cuccaro et al. (2004); Gidney and Ekerå (2021). Moreover, the decoders we consider are readily applicable to other codes and gates, such as the recently developed quantum LDPC codes Breuckmann and Eberhardt (2021), space-time codes Gottesman (2022); Delfosse and Paetznick (2023), and the transversal $S$ gate in the two-dimensional color code, whose decoding problem is challenging for standard decoders Sarvepalli and Raussendorf (2012); Delfosse (2014); Gidney and Jones (2023).
@@
Our results can be extended in several directions. Although the computational runtime of belief-HUF is efficient, its threshold is lower than that of MLE on certain circuits (Fig. 3(c)). Alternative decoders, such as those utilizing tensor networks Bravyi et al. (2014); Ferris and Poulin (2014); Chubb (2021), belief propagation with ordered statistics decoding Panteleev and Kalachev (2021), or machine learning Bausch et al. (2023); Cao et al. (2023); Wang et al. (2023b) could potentially close this performance gap. The runtime of belief-HUF can be further optimized by iteratively decoding fault-tolerant blocks rather than the full circuit Bombín et al. (2023); Skoric et al. (2023); Tan et al. (2023). These techniques can also be adapted to realistic experimental noise models, such as non-Pauli errors, biased noise Cong et al. (2022), and erasure errors Wu et al. (2022); Ma et al. (2023); Sahay et al. (2023). These considerations indicate that by co-designing the logical algorithm, error correcting code, and decoding strategy with experimental hardware in mind, the performance and resource requirements of logical algorithms can continue to be improved Cong et al. (2022); Xu et al. (2023).
@@ 1.2 Belief-HUF decoder
A key component of UF is the termination condition, which signifies that a cluster is satisfied. On the surface code, UF typically has two termination conditions: the total number of flipped checks in that cluster is even, or the cluster contains an error mechanism that flips only one check. Both of these conditions indicate that the cluster is satisfied, since errors in the bulk of the surface code only flip a pair of checks Fowler et al. (2012). Note that the termination condition for a cluster can be evaluated in constant time by tracking the parity of the number of flipped checks, and the existence of an error mechanism in the cluster which flips one check.
@@
$w_{E_{j},C_{i}}=\frac{\log\big{(}\frac{p_{j}}{1-p_{j}}\big{)}}{r_{E_{j}}^{\varepsilon}},$ (S2)
where $r_{E_{j}}$ is the order of the hyperedge $E_{j}$ (meaning the number of vertices it connects in the original decoding hypergraph), and $\varepsilon$ is a hyperparameter of the HUF decoder. In Section 2.2, we compare the performance of HUF for different settings of $\varepsilon.$
@@
The computational runtime of HUF grows polynomially with the size of the decoding Tanner graph $M$. In the worst case, one must merge all vertices into a single cluster to find a satisfiable error configuration, requiring $O(M+N)=O(M)$ total merge operations. By utilizing the union-find data structure Delfosse et al. (2022), the complexity of each merge operation is only $O(\alpha(M+N))$, where $\alpha$ is the slowly-growing inverse Ackermann function. Each merge also requires a satisfaction check (Algorithm 1), which involves solving a linear system over $\mathbb{F}_{2}$. Naively, the linear system can be solved in $O(n^{3})$ time using Gaussian elimination, where $n$ is the number of variables in the system, corresponding to a worst-case total runtime of

### v2

@@ Abstract
Quantum error correction is believed to be essential for scalable quantum computation, but its implementation is challenging due to its considerable space-time overhead. Motivated by recent experiments demonstrating efficient manipulation of logical qubits using transversal gates (Bluvstein et al., Nature 626, 58-65 (2024)), we show that the performance of logical algorithms can be substantially improved by decoding the qubits jointly to account for error propagation during transversal entangling gates. We find that such correlated decoding improves the performance of both Clifford and non-Clifford transversal entangling gates, and explore two decoders offering different computational runtimes and accuracies. In particular, by leveraging the deterministic propagation of stabilizer measurement errors through transversal Clifford gates, we find that correlated decoding enables the number of noisy syndrome extraction rounds between these gates to be reduced from $O(d)$ to $O(1)$ in Clifford circuits, where $d$ is the code distance. We verify numerically that this approach substantially reduces the space-time cost of deep logical Clifford circuits. These results demonstrate that correlated decoding provides a major advantage in early fault-tolerant computation, as realized in recent experiments, and further indicate it has considerable potential to reduce the space-time cost in large-scale logical algorithms.
Quantum error correction (QEC) is believed to be essential for large-scale quantum computation Shor (1995); Preskill (1998); Dennis et al. (2002); Nielsen and Chuang (2010); Kitaev (2003). It is a rapidly developing frontier, with recent milestone demonstrations of the preservation of a logical qubit and universal logic gates Ofek et al. (2016); Reinhold et al. (2020); Ryan-Anderson et al. (2021); Postler et al. (2022); Ryan-Anderson et al. (2022); AI and Collaborators (2023a, b); Honciuc Menendez et al. (2024); Sivak et al. (2023); Wang et al. (2024); Self et al. (2024). However, implementing large-scale algorithms with protected, logical qubits is challenging due to its considerable space-time overhead. Most recently, reconfigurable neutral atom arrays have realized quantum algorithms with tens of logical qubits and hundreds of logical entangling gates Bluvstein et al. (2024), opening the door to early fault-tolerant computation. Key to these results was the ability to efficiently perform fault-tolerant transversal entangling gates by dynamically reconfiguring qubits during the computation Bluvstein et al. (2022), similar to prior work with trapped ions Postler et al. (2022); Ryan-Anderson et al. (2022); Honciuc Menendez et al. (2024); Wang et al. (2024).
@@
Correlated decoding algorithms.— To formalize this approach, we define the decoding hypergraph of the logical algorithm Fowler et al. (2011); Gidney (2021); Gottesman (2022); Higgott et al. (2023); McEwen et al. (2023); Delfosse and Paetznick (2023), which relates each possible physical error mechanism to the stabilizer measurement(s) it flips. The hypergraph vertices correspond to $N$ measured checks of the logical qubits $\{C_{i}\}_{i=1\dots N}$, which compare consecutive stabilizer measurements in time. Each check is defined as the product of a stabilizer measurement and its previous measurement (if one exists), obtained from back-propagating the stabilizer operator through the circuit to the point it was previously measured. If one stabilizer measurement in the check detects an error, the check flips from $+1$ to $-1$. Each error mechanism is associated with a hyperedge connecting the check(s) it flips. Errors copied during transversal entangling gates therefore correspond to hyperedges connecting the checks of multiple logical qubits (Fig. 1(b)). We use Stim Gidney (2021), a Clifford circuit simulator, to identify the $M$ hyperedges $\{E_{j}\}_{j=1\dots M}$ and their probabilities $\{p_{j}\}_{j=1\dots M}$ under a chosen circuit error model, and which error source(s) $I(C_{i})$ flip each check $C_{i}$ (Fig. 1(c)).
@@
Because finding the MLE exactly is NP-hard, we expect the runtime of the above algorithm to grow exponentially in $M$ in the worst-case Berlekamp et al. (1978); Vardy (1997). This motivates our second algorithm, hypergraph union-find with belief propagation (belief-HUF) Delfosse and Nickerson (2021); Delfosse et al. (2022), which instead approximates the MLE and has an efficient worst-case time complexity of $O(M^{3}\log(M))$ sup . Belief-HUF first uses belief propagation to update the error probabilities $\{p_{j}\}_{j=1\dots M}$ to approximate the probability of each error given the measured syndrome Higgott et al. (2023). Leveraging the insight that errors are typically localized, belief-HUF then expands clusters of error mechanisms and checks on the decoding hypergraph at a rate dependent on the error probabilities, merging clusters that meet. A cluster stops expanding when it contains an error mechanism consistent with its measured checks, and the algorithm terminates when all clusters stop expanding.
Decoding transversal Clifford gates.— As a simple illustrative example, we apply these techniques to decode a transversal CNOT, which is available for all Calderbank-Shor-Steane (CSS) codes Shor (1996). A key property of error-correcting circuits is that their performance can be improved for physical error rates below some threshold $p_{\text{th}}$ by increasing the code distance (the minimum weight of a logical Pauli operator), which is related to how many errors can be corrected Dennis et al. (2002); Fowler et al. (2012). In practice, the threshold depends on the decoder, as higher-accuracy decoders can correct more complex patterns of errors. To explore the optimal advantage of correlated decoding, we will compare decoding two surface codes jointly with MLE, and independently with MWPM Higgott (2022); Higgott and Gidney (2023).
@@
Decoding deep logical Clifford circuits.— In realistic quantum algorithms, stabilizer measurements are used to remove entropy from the logical qubit by tracking physical errors over time Gottesman (2009). A challenge, however, is that the stabilizer measurements are also noisy. As a result, entangling gates based on lattice surgery Horsman et al. (2012) and standard error correction constructions Shor (1996); Gottesman (2009) require $d$ rounds of syndrome extraction to ensure fault tolerance to these errors, creating considerable space-time overhead. We now explore if we can perform fewer than $d$ rounds between transversal CNOTs, by leveraging the deterministic propagation of stabilizer measurement errors through CNOTs to verify stabilizer measurements using surrounding rounds of syndrome extraction.
@@
In contrast, because the correlated decoders can be applied to general decoding hypergraphs, they can potentially decode circuits with fewer than $d$ syndrome extraction rounds between CNOTs. To study this, we numerically simulate a deep logical Clifford circuit between four surface codes. The codes are initialized noiselessly in |$\ket{+_{L}}^{\otimes 4}$, then entangled with 32 layers of transversal gates, with $n_{r}$ rounds of syndrome extraction between layers (Fig. 3(b)). Each layer is comprised of a random transversal Hadamard $H_{L}$ or Pauli gate $\{X_{L},Y_{L},Z_{L}\}$, followed by a random pairing of CNOTs with randomized designation of control and target. We add depolarizing noise to each circuit-level physical operation with probability $p$, as in Refs. Higgott et al. (2023); sup . Finally, we noiselessly measure the $X$ and $Z$ stabilizers and the logical stabilizers of the resulting state, then decode to find the logical error rate $P_{L}$ Gottesman (1998). In the Supplementary Materials sup , we confirm that our results are insensitive to boundary effects from the final noiseless measurements, and representative of different randomly sampled circuits.
@@
We first study the case of one round of syndrome extraction per CNOT ($n_{r}=1$). Figure 3(c) shows the logical error rate per round, $P_{L,\text{max}}\big{[}1-(1-P_{L}/P_{L,\text{max}})^{1/(32n_{r})}\big{]}$, where $P_{L,\text{max}}=1-\frac{1}{2^{4}}$ is the error rate of the maximally mixed logical state, as a function of $p$ sup . The numerical results are consistent with the existence of a threshold for both belief-HUF and MLE. Though we cannot reliably fit the thresholds because $P_{L}/P_{L,\text{max}}\to 1$ near the threshold, we estimate that $p_{\text{th}}\geq 0.56\%$ for belief-HUF and $p_{\text{th}}\geq 0.87\%$ for MLE from the largest physical error rate at which $P_{L}$ decreases with $d$ sup . Furthermore, in the Supplementary Materials sup we observe that the MLE logical error rate scaling is exponentially suppressed in $\lfloor\frac{d+1}{2}\rfloor$.
@@
Decoding transversal non-Clifford gates.— Correlated decoding can also be applied to non-Clifford transversal gates. As a proof-of-concept, we study a transversal CCZ between three fifteen-qubit quantum Hamming codes, with perfect syndrome extraction Paetznick and Reichardt (2013); sup . We perfectly initialize each code in an $X$ or $Z$-basis product state, then perform a CCZ with single-physical-qubit depolarizing noise beforehand. Because the CCZ is non-Clifford, these errors can propagate to a superposition of Pauli errors on multiple logical qubits, each generating a hyperedge in the decoding hypergraph (Fig. 5(a-b)) sup ; Bombin (2018); Scruby et al. (2022). We analyze initial states for which the circuit outputs a logical stabilizer state, which can be efficiently simulated sup . We then noiselessly measure the $X$ and $Z$ stabilizers and logical stabilizers, and decode to compute the average logical error rate over initial states Gottesman (1998). Figure 5(c) shows that correlated decoding with MLE outperforms decoding the qubits independently by a factor of $\simeq 1.5$. Note that the gain here is smaller than with Clifford gates, as the error propagation is no longer fully deterministic. This motivates additional analysis of correlated decoding with non-Clifford gates as an avenue for future research.
@@
Outlook.— Our results show that correlated decoding constitutes a powerful tool for improving the performance of logical algorithms. These insights were used in the experiments of Ref. Bluvstein et al. (2024) to entangle a logical Bell pair between surface codes and to fault-tolerantly prepare logical GHZ states. The present work demonstrates that because physical and stabilizer measurement errors can be tracked, these same techniques allow reducing the number of syndrome extraction rounds from $O(d)$ to $O(1)$ in transversal Clifford circuits.
@@
Our correlated decoding technique can be extended in several directions. Although the computational runtime of belief-HUF is efficient, its threshold is lower than that of MLE on certain circuits (Fig. 3(c)). Alternative decoders, such as those utilizing tensor networks Bravyi et al. (2014); Ferris and Poulin (2014); Chubb (2021), belief propagation with ordered statistics decoding Panteleev and Kalachev (2021), or machine learning Torlai and Melko (2017); Liu and Poulin (2019); Varsamopoulos et al. (2020); Kuo and Lai (2022); Bausch et al. (2023); Cao et al. (2023); Maan and Paler (2024); Wang et al. (2023); Ninkovic et al. (2024) could potentially close this performance gap. The runtime of belief-HUF can be further optimized by iteratively decoding fault-tolerant blocks rather than the full circuit Bombín et al. (2023); Skoric et al. (2023); Tan et al. (2023). Moreover, the decoders we consider are readily applicable to other codes and gates, such as the recently developed high-rate quantum LDPC codes Tillich and Zémor (2014); Panteleev and Kalachev (2021); Leverrier and Zémor (2022); Panteleev and Kalachev (2022a, b); Breuckmann and Eberhardt (2021); Xu et al. (2024), space-time codes Gottesman (2022); Delfosse and Paetznick (2023), and the transversal $S$ gate in the two-dimensional color code Sarvepalli and Raussendorf (2012); Delfosse (2014); Gidney and Jones (2023). These techniques can also be adapted to realistic experimental noise models, such as non-Pauli errors, biased noise Cong et al. (2022), and erasure errors Wu et al. (2022); Ma et al. (2023); Sahay et al. (2023).
Another important question involves the extent to which the present techniques can be used to achieve practical speedups in classically complex circuits involving non-Clifford operations and feed-forward. In a recent complementary manuscript, we show that the methods described in this work can indeed be utilized for an $O(d)$ time reduction for such universal computation Zhou et al. (2024). Taken together, these results demonstrate that correlated decoding is emerging as a core building block both in new theories of fault-tolerance and in practical reductions to the cost of large-scale computation Zhou et al. (2024). Improving the decoders and extending this technique to new applications are thus key next steps.
@@ 1.2 Belief-HUF decoder
A key component of UF is the termination condition, which signifies that a cluster is satisfied. On the surface code, UF typically has two termination conditions: the total number of flipped checks in that cluster is even, or the cluster contains an error mechanism that flips only one check. Both of these conditions indicate that the cluster is satisfied, since errors in the bulk of the surface code only flip a pair of checks Fowler et al. (2012). Note that this termination condition for a cluster can be evaluated in constant time by tracking the parity of the number of flipped checks, and the existence of an error mechanism in the cluster which flips one check.
@@
$w_{E_{j},C_{i}}=\log\Big{(}\frac{p_{j}}{1-p_{j}}\Big{)}r_{E_{j}}^{\varepsilon},$ (S2)
where $r_{E_{j}}$ is the order of the hyperedge $E_{j}$ (meaning the number of vertices it connects in the original decoding hypergraph), and $p_{j}$ is the probability that error $E_{j}$ occured. $\varepsilon$ is a hyperparameter of the HUF decoder controlling how quickly clusters expand along higher-order hyperedges, which we take to be either $-1$ or 00. In Section 2.2, we compare the performance of HUF for different settings of $\varepsilon.$
@@
The computational runtime of HUF grows polynomially with the size of the decoding Tanner graph $M$. In the worst case, one must merge all vertices into a single cluster to find a satisfiable error configuration, requiring $O(M+N)=O(M)$ total merge operations. By utilizing the union-find data structure, the complexity of each merge operation is only $O(\alpha(M+N))$, where $\alpha$ is the slowly-growing inverse Ackermann function (for all practical purposes, $\alpha\leq 3$Delfosse et al. (2022). Each merge also requires a satisfaction check (Algorithm 1), which involves solving a linear system over $\mathbb{F}_{2}$. Naively, the linear system can be solved in $O(n^{3})$ time using Gaussian elimination, where $n$ is the number of variables in the system, corresponding to a worst-case total runtime of